OpenStudy (enricoanders):

Jane has a pre-paid cell phone with Splint. She can't remember the exact costs, but her plan has a monthly fee and a charge for each minute of calling time. In June she used 260 minutes and the cost was $49.00. In July she used 940 minutes and the cost was $117.00. A) Express the monthly cost C as a function of x, the number of minutes of calling time she used. B) If Jane used 504 minutes of calling time in August, how much was her bill?

OpenStudy (enricoanders):

@DanJS

OpenStudy (enricoanders):

@sshayer

OpenStudy (danjs):

The total cost is a linear equation, and they give you 2 points of data. You can make a line equation.

OpenStudy (danjs):

Total Cost = Fee + Minute Charge * #minutes The two points given are, (x , C) = (260 , 49.00) (x1 , C1) = (940, 117.00) Remember the point slope line equation y - y1 = m*(x - x1) C - C1 = m*(x - x1) Use the two points for that

OpenStudy (sshayer):

x+260y=49 x+910y=117 find x and y

OpenStudy (enricoanders):

I tried that and it was wrong ill try again

OpenStudy (sshayer):

not 910 but 940

OpenStudy (enricoanders):

y=1/10x+23

OpenStudy (enricoanders):

is that correct?

OpenStudy (enricoanders):

@sshayer @DanJS

OpenStudy (sshayer):

correct.

OpenStudy (enricoanders):

can you help me with part b

OpenStudy (sshayer):

use x=504

OpenStudy (enricoanders):

so i do y-49=1/10(x-504)

OpenStudy (sshayer):

\[C=\frac{ 1 }{ 10 }\times 504+23=?\]

OpenStudy (enricoanders):

omg im so off haha let me do that

OpenStudy (enricoanders):

527/10

OpenStudy (sshayer):

50.40+23.00=

OpenStudy (enricoanders):

73.4

OpenStudy (sshayer):

correct

OpenStudy (enricoanders):

sp thats the answer?

OpenStudy (enricoanders):

can you explain what you did there please? i have a test tomorrow and i dont quite understand 100%

OpenStudy (sshayer):

in your equation x is the number of minutes

OpenStudy (danjs):

(x , C) = (260 , 49.00) (x1 , C1) = (940, 117.00) Those two points give a slope m=1/10 Using that slope m=1/10 and the point (260 , 49.00) C - 49 = (1/10)*(x - 260) C - 49 = (1/10)*x - 26 C = (1/10)*x + 23 The function is C(x) = (1/10)*x + 23

OpenStudy (enricoanders):

I get that part, i dont get the second part

OpenStudy (danjs):

For x is 504 minutes... C(504) = (1/10)*(504) + 23 = 367/5 = 73.40 dollars

OpenStudy (enricoanders):

ohh okayy! thank you so much!!